OpenStudy (anonymous):

A gun is fired horizontally at the center of a target's bull's eye. If the target is located 100m away and the bullet is traveling 500 m/s, how far beneath the center of the bull's eye does the bullet impact?

OpenStudy (4n1m0s1ty):

hint: Use these equations \[y=-0.5g t^{2}\] \[x=v_{0}t\]

OpenStudy (anonymous):

100/500=.2s y=-.5*g*t^2= -.5*9.8*(.2^2)=-.196 m